Reports until 14:47, Wednesday 21 August 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 14:47, Wednesday 21 August 2024 - last comment - 13:09, Thursday 29 August 2024(11845)
OFC1 Super optic cleaning

Super optic cleaning prep notes:

I've moved the bin of extra OFC parts to the other table in order to clear room on the OFC table. This will allow us to keep the exposed super optics inside the clean room at all times. I've set up a prep area on the OFC table closest to the door to B102. I have cleaned a pair of metal tweezers (the plastic ones in trials were very staticy), all of the required hex keys, and moved the first contact supplies to this location.

Currently my plan is to fully remove the ~2 inch spacer/flexture mount combination along with the optic. It would be possible to unscrew the nylon tip set screw that holds the mirrors in place but this seems like it adds unneeded risk and the flexture mount combinations can be removed/replaced and still retain inter-cavity alignment.

The piezo mirror must be fully removed in order to access the mirror. I have set aside an SM1 ring spanner wrench and piezo alignment tool to be cleaned when I am garbed up.

Two of the mirrors on one side (OFC1:FCM2 and OFC1:FCM4) are too close to a beamsplitter (OFC1:BS2) to remove. This BS must be removed from the table temporarily. I have done this.

End of preparation for first contacting particle count 0/0/0 and 1/0/0 during two 60 second count trials.

** Lunch**

Midway particle count: 0/0/0

Plan is to clean both the HR and AR sides. I think the way to do it is rest the optic on the AR side, paint the HR side, let this side dry, flip it and paint the AR side while the optic is resting on the painted HR side, wait til both are dry, remove both sides without resting the optic and install back into mount.

HR side first contact complete. Piezo mirror was visibly dirty. Looked like metal shavings. The AR side also had a large imprint on it from the piezo being driven a couple of times. Similar to what we've seen in the past.

AR side first contact complete. Mirrors are resting on dry HR side. Waiting 10 more minutes and then peeling.

Cleaned the inside circular part of the spacers that hold the mirror. The piezo mirror mount one had what looked like small metalic particulates.

All 8 peels were very successful. I noticed some specs (most likely stringing from the FC) on the AR side of OFC1:FCM1. Repainted this side while it was in its mount. Peel was successful. Reinstalled the mirrors on the cavity and now will scan and look to realign.

End of work particle count: 0/0/0

 

Comments related to this report
Torrey Cullen - 13:09, Thursday 29 August 2024 (11850)

Forgot to provide photos taken during the cleaning. See attached.

Images attached to this comment
Torrey Cullen - 15:15, Thursday 22 August 2024 (11846)

It seems like the alignment is not recoverable. Tried a few tricks to quickly regain alignment but it has not worked. Will need to start from scratch.